# PRING: Rethinking Protein-Protein Interaction Prediction from Pairs to Graphs

### v1

@@ D.2.3 Essential Protein Justification
[…]
@@
[…]
@@ Appendix F Experimental Settings
Specifically, PLM-interact (35M and 650M variants) is used in inference mode with publicly released pre-trained weights. Due to computational constraints, RF2-Lite is evaluated on only 10% of the test graphs (a graph is randomly sampled for each node size to evaluation). All other baseline models are trained from scratch. In addition, we conduct a case study of Chai-1 on just three Human graphs (node sizes range from 20 to 60).
For the SPRINT, which does not need GPUs for inference, the experiments are done on Kunpeng-920, a 64-core platform. For the model trained from scratch, NVIDIA A100-PCIE-40GB is used with the same CPU support. Additionally, for the RF2-Lite and Chai-1, the experiments are conducted on NVIDIA-A800-SXM4-80GB. Since RF2-Lite requires MSAs as input, we use Uniref90 [97] as the search database with GPU-accelerated version of MMseq2 [105]. Given that Chai-1 performs well without the use of MSAs or templates, and that generating MSAs is computationally intensive, we opt to use single-sequence inputs exclusively.
@@ Appendix G Experimental Results
We present detailed results for all evaluated models across both topology-oriented and function-oriented tasks. The following sections offer a comprehensive comparison of model performance, highlighting their ability to preserve structural properties and capture biologically meaningful patterns.
@@
Additionally, we include a case study on Chia-1 [99], a state-of-the-art protein structure prediction model, to evaluate its potential in PPI prediction.
Finally, we conduct a scaling-law analysis to investigate how model size correlates with graph-level performance, offering insights into the trade-offs between model complexity and predictive fidelity.
@@ G.1 Topology-oriented Task
Naive Sequence Topsy-Turvy 0.183 1.74 17.3 11.4 14.6 0.130 3.22 180 201 65.9 0.232 4.88 104 64.8 36.7 5
@@
Naive Sequence Topsy-Turvy 0.508 0.596 0.120 0.160 0.541 9
@@ G.1.1 Intra-species PPI Network Construction
Naive Sequence Topsy-Turvy 0.163 1.91 20.6 20.5 15.6 0.153 2.32 13.0 123 61.3 0.267 3.90 60.5 33.0 30.6 4
@@
Naive Sequene Topsy-Turvy 0.501 0.696 0.118 0.158 0.556 10
@@
Naive Sequence Topsy-Turvy 0.142 1.55 11.6 14.2 10.9 0.130 2.58 115 101 28.5 0.186 5.12 100.0 50.1 35.0 4
@@
Naive Sequene Topsy-Turvy 0.509 0.529 0.168 0.255 0.506 10
@@
Naive Sequence Topsy-Turvy 0.220 5.54 132 43.6 39.0 0.0810 3.26 312 126 50.1 0.160 10.8 422 152 70.4 8
@@
Naive Sequene Topsy-Turvy 0.507 0.538 0.211 0.219 0.539 9
@@ G.1.3 Correlation Between Binary Classification Scores and Network Metrics
To investigate whether standard classification metrics reflect the structural quality of reconstructed PPI networks, we analyze their correlation with five topology-aware metrics across all evaluated models. As shown in Fig. 7, classification metrics such as recall, F1 score, and AUPR exhibit strong positive correlations with graph-level metrics like graph similarity (e.g., Pearson $r=0.901$ between AUPR and graph similarity). In contrast, accuracy and precision show weak correlations with certain topology metrics, particularly clustering coefficient and spectral distance, suggesting that high pairwise prediction scores may not always align with biologically meaningful network structures. These results underscore the importance of evaluating PPI prediction models beyond binary classification, as traditional metrics alone may fail to reflect structural coherence and biological relevance.
@@ G.2 Function-oriented Task
Naive Sequene Topsy-Turvy 0.240 0.613 0.338 9
@@ G.2.1 Protein Complex Pathway Prediction
Category GO:BP GO:MF GO:CC Average
[…]
@@
Naive Sequene Topsy-Turvy 0.131 0.333 0.242 0.652 0.323 0.919 0.232 0.635 7
@@ G.2.2 GO Enrichment Analysis
Naive Sequence Topsy-Turvy 0.43 0.755 7
@@ G.2.3 Essential Protein Justification
G.3 Case Study

### v2

@@ D.2.3 Essential Protein Justification
$\text{Precision@K}=\frac{|\text{Top}_{K}\cap\text{Essential}|}{K}$ (12)
@@
$\text{Distribution Overlap}=\int_{-\infty}^{\infty}\min\left(p_{\text{essential}}(x),\,p_{\text{non-essential}}(x)\right)dx$ (13)
@@ Appendix F Experimental Settings
Specifically, PLM-interact (35M and 650M variants) is used in inference mode with publicly released pre-trained weights. Due to computational constraints, RF2-Lite is evaluated on only 10% of the test graphs (a graph is randomly sampled for each node size for evaluation). All other baseline models are trained from scratch. In addition, we conduct a case study of Chai-1 on just three Human graphs (node sizes range from 20 to 60). For the binary-based PPI prediction models, we use the default probability threshold of 0.5 to distinguish between interacting and non-interacting pairs.
The experiments are conducted on heterogeneous hardware platforms. For the SPRINT, which does not need GPUs for inference, the experiments are done on Kunpeng-920, a 64-core platform. For the model trained from scratch, NVIDIA A100-PCIE-40GB is used with the same CPU support. Additionally, for the RF2-Lite and Chai-1, the experiments are conducted on NVIDIA-A800-SXM4-80GB. Since RF2-Lite requires MSAs as input, we use Uniref90 [105] as the search database with GPU-accelerated version of MMseq2 [119]. Given that Chai-1 performs well without the use of MSAs or templates, and that generating MSAs is computationally intensive, we opt to use single-sequence inputs exclusively.
@@ Appendix G Experimental Results
We present detailed results for all evaluated models across both topology-oriented (Sec. G.1) and function-oriented tasks (Sec. G.2). The following sections offer a comprehensive comparison of model performance, highlighting their ability to preserve structural properties and capture biologically meaningful patterns.
@@
We further include the complete test results from the ablation study on varying positive-to-negative PPI training ratios in Sec. G.3. We analyze the impact of different probability thresholds on the trade-off between precision and recall in Sec. G.4 as well.
Additionally, we include a case study on Chai-1 [107] in Sec. G.5, a state-of-the-art protein structure prediction model, to evaluate its potential in PPI network construction.
Finally, we conduct a scaling-law analysis in Sec. G.6 to investigate how model size correlates with graph-level performance, offering insights into the trade-offs between model complexity and predictive accuracy.
@@ G.1 Topology-oriented Task
Naive
Sequence
Topsy-Turvy 0.183 1.74 17.3 11.4 14.6 0.130 3.22 180 201 65.9 0.232 4.88 104 64.8 36.7 5
@@
Naive
Sequence
Topsy-Turvy 0.508 0.596 0.120 0.160 0.541 9
@@ G.1.1 Intra-species PPI Network Construction
Naive
Sequence
Topsy-Turvy 0.163 1.91 20.6 20.5 15.6 0.153 2.32 13.0 123 61.3 0.267 3.90 60.5 33.0 30.6 4
@@
Naive
Sequene
Topsy-Turvy 0.501 0.696 0.118 0.158 0.556 10
@@
Naive
Sequence
Topsy-Turvy 0.142 1.55 11.6 14.2 10.9 0.130 2.58 115 101 28.5 0.186 5.12 100.0 50.1 35.0 4
@@
Naive
Sequene
Topsy-Turvy 0.509 0.529 0.168 0.255 0.506 10
@@
Naive
Sequence
Topsy-Turvy 0.220 5.54 132 43.6 39.0 0.0810 3.26 312 126 50.1 0.160 10.8 422 152 70.4 8
@@
Naive
Sequene
Topsy-Turvy 0.507 0.538 0.211 0.219 0.539 9
@@ G.1.2 Cross-species PPI network Construction
We further account for numerical phylogenetic distance to provide a deeper understanding of cross-species generalization. To investigate this, we conducted an ablation study by computing the Pearson correlation between model performance and the estimated phylogenetic distance from Human. Specifically, we set the intra-species (Human→Human) distance to 0, and adopt widely accepted phylogenetic estimates [120] for other species: Human→Arath (1.5 Gyr), Human→Yeast (1.2 Gyr), and Human→Ecoli (3.6 Gyr), which consist the distance vector $\mathbf{d}=[0,1.5,1.3,1.6]$. For given model $M$ and performance metric (e.g., graph similarity (GS)), let $\mathbf{p}_{m}^{M}=[p_{\rm{Human}},p_{\rm{Arath}},p_{\rm{Yeast}},p_{\rm{Ecoli}}]$ denotes the model’s performance vector across the four species. We calculate the Pearson correlation $r(\mathbf{d},\mathbf{p}_{m}^{M})$ to quantify the relationship between phylogenetic distance and model performance. This analysis is conducted on four representative baselines, and are summarized in Tab. 15.
Table 15: Pearson correlation between phylogenetic distance and graph-level metrics.
Sampling Method BFS DFS Random Walk
Category Method $r(\mathbf{d},\mathbf{p}_{\rm{GS}})$ $r(\mathbf{d},\mathbf{p}_{\rm{RD}})$ $r(\mathbf{d},\mathbf{p}_{\rm{Deg.}})$ $r(\mathbf{d},\mathbf{p}_{\rm{Clus.}})$ $r(\mathbf{d},\mathbf{p}_{\rm{Spectral}})$ $r(\mathbf{d},\mathbf{p}_{\rm{GS}})$ $r(\mathbf{d},\mathbf{p}_{\rm{RD}})$ $r(\mathbf{d},\mathbf{p}_{\rm{Deg.}})$ $r(\mathbf{d},\mathbf{p}_{\rm{Clus.}})$ $r(\mathbf{d},\mathbf{p}_{\rm{Spectral}})$ $r(\mathbf{d},\mathbf{p}_{\rm{GS}})$ $r(\mathbf{d},\mathbf{p}_{\rm{RD}})$ $r(\mathbf{d},\mathbf{p}_{\rm{Deg.}})$ $r(\mathbf{d},\mathbf{p}_{\rm{Clus.}})$ $r(\mathbf{d},\mathbf{p}_{\rm{Spectral}})$
Naive Seq. D‑SCRIPT ‑0.82 0.94 0.61 0.92 0.93 ‑0.70 0.81 0.65 0.63 ‑0.30 ‑0.69 0.91 0.89 0.81 0.93
PPITrans ‑0.98 0.80 0.94 0.56 1.00 ‑0.90 0.89 0.53 0.81 0.79 ‑0.91 0.86 0.87 0.77 0.95
PLM TUnA ‑0.97 0.97 0.96 0.94 0.92 ‑0.98 0.98 0.49 0.70 0.64 ‑0.96 0.95 0.92 0.86 0.97
Structure Struct2Graph ‑0.98 ‑0.26 1.00 ‑0.01 ‑0.55 ‑0.85 ‑0.03 0.43 0.98 ‑0.53 ‑0.88 0.63 0.84 0.17 0.61
We find that phylogenetic distance is strongly correlated with model performance degradation across species. For all models and sampling strategies, GS consistently exhibits a strong negative correlation with phylogenetic distance (correlations typically < -0.7). This indicates that as phylogenetic distance increases, GS decreases—i.e., model predictions become less similar to the ground-truth network. Most other metrics show strong positive correlations with phylogenetic distance. This means that as the evolutionary gap widens, predicted networks increasingly diverge from true biological networks, reflected by higher RD, degree, clustering, or spectral values, all of which signal degraded performance.
Table 16: Alternative Cross-species Transfer.
Transfer Direction Arath→Arath Arath→Yeast Arath→Ecoli
Category Method GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$ GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$ GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$
Naive Seq. D‑SCRIPT 0.407 2.24 38.4 11.6 19.0 0.288 3.68 52.3 25.0 28.9 0.243 6.11 121 46.1 44.1
PPITrans 0.470 2.96 43.7 31.4 17.8 0.376 4.68 78.0 50.8 35.8 0.312 5.14 138 51.4 44.1
PLM TUnA 0.548 1.54 22.6 19.1 14.4 0.378 3.87 58.5 42.9 27.0 0.302 5.39 129 49.9 42.1
Structure Struct2Graph 0.352 2.59 67.0 70.5 16.3 0.266 4.84 92.9 62.1 30.6 0.218 5.53 145 61 38.8
In addition to the cross-species transfer direction presented in the main manuscript, we perform new experiments on botany-to-fungi (Arath → Yeast) and botany-to-bacteria (Arath → Ecoli) transfer learning. We followed the same data construction pipeline described in Sec. 3.1 to avoid data leakage. We focused on the BFS sampling strategy and evaluated four representative models. The results are presented in Tab. 16, and they confirm our previous experimental observation: cross-species generalization becomes more challenging with greater evolutionary divergence.
Specifically, compared to the intra-species setting (Arath → Arath), the GS scores for Arath → Yeast and Arath → Ecoli decreased by 26% and 39%, respectively. Other graph-level metrics observe performance degradation as well, indicating the performance of the transfer learning decreases with the increased evolutionary distances.
@@ G.1.3 Correlation Between Binary Classification Scores and Network Metrics
To investigate whether standard classification metrics reflect the structural quality of reconstructed PPI networks, we analyze their correlation with five topology-aware metrics across all evaluated models. We summarize some important discoveries in Fig. 8:
  • Recall often distorts the network topology: Recall measures the proportion of true positives correctly identified. To achieve higher recall, the model typically predicts more edges to capture as many true interactions as possible, even at the cost of including false positives. This leads to a denser predicted graph, which deviates from the sparsity characteristic of real PPI networks, ultimately hurting the preservation of the original topological structure.
  • Precision tends to preserve network topology: Precision measures the proportion of correct positive predictions. Improving precision requires minimizing false positives, which makes the model more conservative in predicting edges. As a result, the predicted graph becomes sparser and more aligned with the intrinsic sparsity of real PPI networks, thereby helping to preserve the underlying network topology.
  • Composite metrics like F1-score and AUPR may hide topological differences: Since these metrics balance both precision and recall, they may yield high scores even when the predicted network substantially deviates from the true topology. As a result, their correlation with topological fidelity tends to be weak, potentially masking structural distortions in the predicted network.
These results underscore the importance of evaluating PPI prediction models beyond binary classification, as traditional metrics alone may fail to reflect structural coherence and biological relevance.
@@ G.1.4 Impact of Subgraph Size on Topology-Oriented Performance
Naive
Sequene
Topsy-Turvy 0.240 0.613 0.338 9
@@
Category GO:BP GO:MF GO:$C$ Average
Category Model $uparrow FA\uparrow CR\uparrow$ $uparrow FA\uparrow CR\uparrow$ Avg. Rank
@@
Naive
Sequene
Topsy-Turvy 0.131 0.333 0.242 0.652 0.323 0.919 0.232 0.635 7
@@
G.2 Function-oriented Task
G.2.1 Protein Complex Pathway Prediction
@@ G.2.2 GO Enrichment Analysis
Naive
Sequence
Topsy-Turvy 0.43 0.755 7
@@ G.2.3 Essential Protein Justification
Table 20: Ablation results on the positive-to-negative ratio of training pairs under three sampling configurations.
BFS
Sampling Method Ratio 1:1 Ratio 1:5 Ratio 1:10
Category Model GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$ GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$ GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$
Naive
Sequence
D-SCRIPT 0.215 1.13 15.4 18.8 29.7 0.032 0.0705 94.2 46.6 83.2 0.035 0.0400 116 54.1 97.2
PPITrans 0.362 3.39 52.2 29.1 26.3 0.393 1.17 8.12 7.37 10.9 0.374 0.692 11.3 7.23 16.4
PLM TUnA 0.342 2.99 47.8 30.2 22.9 0.346 1.14 10.7 9.06 8.77 0.314 0.579 14.4 9.58 17.3
Structure Struct2Graph 0.270 5.58 95.1 50.5 39.2 0.236 4.11 62.6 35.2 25.6 0 5.71E‑05 146 65.8 117
DFS
Sampling Method Ratio 1:1 Ratio 1:5 Ratio 1:10
Category Model GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$ GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$ GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$
Naive
Sequence
D-SCRIPT 0.123 6.40 302 237 67.0 0.0649 0.0545 518 137 136 0.089 0.27 413 112 107
PPITrans 0.314 4.54 418 270 46.6 0.433 2.19 213 153 28.3 0.520 1.19 100 40.0 38.0
PLM TUnA 0.289 4.44 416 272 46.0 0.417 1.49 127 85.6 19.7 0.414 0.933 114 21.9 38.9
Structure Struct2Graph 0.132 10.8 508 321 67.6 0.118 7.29 430 282 50.9 1.74E‑05 3.50E‑05 593 154 145
Random Walk
Sampling Method Ratio 1:1 Ratio 1:5 Ratio 1:10
Category Model GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$ GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$ GS $\uparrow$ RD $\rightarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Spectral $\downarrow$
Naive
Sequence
D-SCRIPT 0.209 5.82 134 64.8 52.0 0.135 0.334 83.2 51.9 55.6 0.01 0.01 188 133 134
PPITrans 0.449 2.57 71.9 46.1 19.6 0.564 1.27 10.8 10.8 12.3 0.547 0.808 23.7 14.9 25
PLM TUnA 0.450 2.23 55.4 42.5 13.7 0.463 0.764 15.8 9.58 13.2 0.425 0.539 38.0 27.6 29.7
Structure Struct2Graph 0.237 5.83 155 103 38.3 0.216 3.44 87.7 72.0 18.4 0 5.71E‑05 197 135 140
G.3 Ablation Study on Positive-to-Negative Ratio
To examine the effect of class imbalance on PPI network reconstruction, we vary the positive-to-negative ratio of training pairs from 1:1 to 1:10 under three sampling configurations: BFS, DFS, and Random Walk. This setting reflects the inherent sparsity of real PPI networks, where non-interacting protein pairs largely outnumber interacting ones. We evaluate representative sequence-, PLM-, and structure-based models using five network-level metrics.
Tab. 20 reports the complete results. Across sampling strategies, moderate imbalance (e.g., 1:5) tends to improve the robustness of PLM-based models (PPITrans, TUnA), while excessive imbalance (1:10) causes performance degradation and overly sparse reconstructions. In contrast, simpler models such as D-SCRIPT and Struct2Graph deteriorate rapidly as the ratio increases, indicating their sensitivity to biased supervision.
Table 21: Effect of Probability Thresholds on Precision and Recall.
PIPR D‑SCRIPT PPITrans Struct2Graph
Threshold Precision Recall Precision Recall Precision Recall Precision Recall
0.2 0.552 0.634 0.570 0.135 0.627 0.821 0.502 0.937
0.3 0.554 0.615 0.582 0.115 0.651 0.769 0.504 0.900
0.5 0.557 0.581 0.597 0.104 0.658 0.738 0.511 0.696
0.7 0.561 0.550 0.764 0.067 0.753 0.535 0.532 0.007
0.8 0.562 0.530 0.780 0.053 0.798 0.438 0.532 0.007
G.4 Ablation Study on Probability Threshold
We evaluate how varying probability thresholds influence the balance between precision and recall across four representative models. As shown in Tab. 21, Struct2Graph and D-SCRIPT are highly sensitive to thresholding, with recall dropping sharply, indicating that a large portion of their prediction scores have low confidence and are clustered around 0.5. In contrast, PIPR and PPITrans exhibit more stable trends: recall gradually decreases while precision increases as the threshold rises, suggesting a trade-off between the two metrics and a limited ability to maintain a strong balance.
G.5 Case Study